OpenStudy (123qwet12):

How would I find the solutions to the equation |4x-6|-1=9?

OpenStudy (anonymous):

| | Means the signs of the numbers are to be changed

OpenStudy (123qwet12):

yea the absolute value signs

OpenStudy (anonymous):

-4x+6-1=9

OpenStudy (123qwet12):

I initially solved the problem and got x=1 and x= 1/2 but those were incorrect

OpenStudy (123qwet12):

alright

OpenStudy (anonymous):

-4x+6-1=9 solve for x

OpenStudy (123qwet12):

the equation this time got me x=-1 and that's one of the correct answers, so I need to solve for the other

OpenStudy (anonymous):

That is correct

OpenStudy (123qwet12):

the second one got me x=1 which is incorrect

OpenStudy (mathmale):

Daniel's approach is fine for half the problem, but there's another half. How would I find the solutions to the equation |4x-6|-1=9? this equation can and should be broken up into two parts: Part 1 is Daniel's: +(4x-6) - 1 = 9 Part 2 is -(4x-7) -1 = 9

OpenStudy (mathmale):

Solve each of these two linear equations, separately. Check each answer in the original equation.

OpenStudy (123qwet12):

I'm talking about the second part of the equation

OpenStudy (123qwet12):

@mathmale I still got an incorrect answer

OpenStudy (123qwet12):

it says one of the answers should be x=4, but I'm unsure how to get that answer

OpenStudy (anonymous):

what was your second equation

OpenStudy (123qwet12):

well to solve this type of equation, you have to split it into two parts I'm not sure how I have to split it, but you gave me the first type, -4x+6-1=9

OpenStudy (anonymous):

i think x is 1

OpenStudy (123qwet12):

I found out how to do it. 1. You have to isolate the absolute value sign 2. THEN you create the two problems 3. solve the 2 problems 4. then substitute to ensure that you have no extraneous solutions

OpenStudy (123qwet12):

thank you for your help @Daniel56k and @mathmale

OpenStudy (anonymous):

can i see the solution

OpenStudy (123qwet12):

|4x-6|-1=9 add 1 to both sides to isolate the absolute value equation |4x-6|=10 then remove the signs 4x-6=10 solve this equation first 4x=16 Divide x=4 4x-6=-10 Create the next problem 4x=-4 Add six to both sides x=-1 Divide

OpenStudy (anonymous):

Wow that is brilliant

OpenStudy (mathmale):

Thanks for your perseverance!